# Curriculum-Guided Layer Scaling for Language Model Pretraining

### v1

@@ 3.1 Data Curricula
Then, by choosing mixture weights over the stratified splits of DCLM, our data schedule (see Table 2) reflects a gradual shift from general to specialized content as model depth increases.
3.2 Compute-Matched Method Comparisons
In our experiments at the GPT-2-Small and LLaMA-3.2-1B parameter counts, we fix a compute budget in pretraining FLOPs across all methods to enable clean comparisons. While prior work on LLM scaling suggests that optimal training lies near the “Chinchilla” compute/data tradeoff curve [23], we select a lower FLOP budget due to academic compute constraints, using 700M tokens for the GPT-2-small scale experiments and 2.5B tokens for the Llama-3.2-1B scale experiments. Following well-established practices in the LLM scaling laws literature [26], we approximate the total compute budget for each experiment, in FLOPs, as:
@@ 3.2 Compute-Matched Method Comparisons
CGLS Compute-Matched Configurations.
Our experiments consisted of training several compute-matched models under our CGLS framework, with varying dataset splits and starting number of layers $N_{1}$. These are summarized as follows, with detailed data splits shown in Tables 1 and 2.
  • CGLS ($N_{1}=C)$: These configurations follow a shared methodology across both model scales. Training begins with $C$ layers trained on a curriculum-balanced dataset, followed by a staged progression to the full model size. At each stage, newly added layers are first trained with a fixed data split (used for all initialization phases), before the full model is fine-tuned on a progressively more challenging data distribution.
Compute-Matched Baselines.
  • Curricularized baseline: This baseline simulates a coarse curriculum by training the full model sequentially on increasingly complex data subsets. For GPT-2-Small, we train on TinyStories, then BookCorpus, then DCLM. For LLaMA-3.2-1B, we follow the same staged training but used the stratified DCLM splits (high school, undergraduate, graduate).
  • MIDAS: MIDAS progressively stacks layers across training stages but does not control the complexity of the training data. We reimplemented the method from pseudocode provided in the original paper [39], as there is no public codebase, with PROP-3 stage budgets, a block size of 3 for GPT-2-Small, and a block size of 4 for LLaMA-3.2-1B.
Further CGLS Implementation Details.
At GPT-2-Small parameter scale, we scaled from 6 to 12 layers in 3 stages; at Llama-3.2-1B parameter scale, from 8 to 16 layers in 4 stages. The train FLOP budgets for each stage were optimized and allocated to favor later, larger stages (20% each for the first three stages; 40% for the final stage at Llama-3.2-1B scale), ensuring sufficient resources for fine-tuning on the most complex data. Additionally, we allocated 20% of the FLOPs for each stage to training only the new layers, and 80% for fine-tuning the full-unfrozen model. Additional details about our experimental setup can be found in A.3.
3.3 Evaluation
To measure generalization and reasoning capabilities, we evaluated the models on several downstream tasks:
  • PIQA (Physical Interaction QA) [5]: This dataset tests physical commonsense reasoning by presenting multiple-choice questions about everyday scenarios.
  • ARC Easy and Challenge [8]: The AI2 Reasoning Challenge benchmarks measure a model’s ability to answer elementary and middle-school science questions. The "Easy" subset contains relatively straightforward questions, while the "Challenge" subset consists of harder, multi-step reasoning tasks.
  • HellaSwag [50]: This task evaluates common-sense reasoning and next-word prediction, requiring the model to identify the most plausible continuation for a given context.
  • MMLU (Massive Multitask Language Understanding) [22]: MMLU consists of multiple-choice questions across 57 tasks spanning the humanities, STEM, social sciences, and more. Given our constrained academic budget for compute, raw accuracy on MMLU and other multiple-choice tasks may not reflect a meaningful signal. Instead, we evaluate Pointwise Mutual Information (PMI) Accuracy and Token-Normalized Accuracy [32, 19], metrics that better account for length and bias effects in generative modeling to obtain a low-variance signal at small scales. These metrics are defined as:
    $\text{acc}_{\text{pmi}}=\arg\max_{i}\ln\frac{P(a_{i}\mid q)}{P(a_{i}\mid u)},\quad\text{acc}_{\text{token}}=\arg\max_{i}\frac{\ln P(a_{i}\mid q)}{\texttt{num\_tokens}(a_{i})},$
@@ 3.3 Evaluation
  • LAMBADA [36]: An open-book language modeling task designed to assess a model’s ability to perform long-range dependency reasoning. We report zero-shot accuracy, computed as whether the model correctly predicts the final word of a passage.
All tasks are evaluated in a zero-shot setting to measure base model quality without additional fine-tuning. We report the token-normalized accuracy, as described above, for all multiple-choice tasks, and PMI for MMLU. We additionally compute perplexity on a 180,000 sample subset of the Pile dataset, capturing how well the methods model diverse, noisy data from a variety of domains.
4 Results
Our experiments evaluate the performance of the CGLS training strategy at two parameter counts (matched to GPT-2-Small and Llama-3.2-1B) across perplexity and zero-shot accuracy on several downstream tasks. Table 3 summarizes the findings, comparing CGLS to several baselines with matched compute budgets in FLOPs. Across both model scales, CGLS consistently outperforms baseline approaches on reasoning-focused benchmarks such as ARC and PIQA, demonstrating its effectiveness in improving generalization under fixed compute budgets. CGLS shows particular strength in structured and knowledge-intensive tasks, supporting the benefits of aligning model growth with data complexity.
Model
Pile
Perplexity
PIQA ARC-E ARC-C HellaSwag Lambada
MMLU
Average
GPT-2-Small
Pretrained GPT-2-Small 23.97 62.51% 39.48% 22.70% 31.14% 38.96%
Baseline (Randomized) 37.51 58.00% 36.11% 21.08% 26.60% 35.45%
Baseline (Curricularized) 36.09 59.09% 36.78% 22.18% 26.43% 36.12%
Layer Scaling Only 39.67 58.76% 35.82% 21.42% 26.35% 35.59%
MIDAS [39] 49.90 56.91% 34.81% 21.16% 25.66% 34.64%
CGLS (Simple) 35.93 59.03% 37.67% 21.76% 19.50% 34.49%
CGLS ($N_{1}=3$) 35.48 59.19% 37.46% 21.08% 26.77% 36.13%
CGLS ($N_{1}=6$) 34.33 60.12% 36.62% 22.53% 26.55% 36.46%
Llama-3.2-1B
Pretrained Llama-3.2-1B 9.25 74.32% 60.56% 36.52% 63.67% 62.00% 31.50% 54.76%
Baseline (Randomized) 20.83 59.09% 36.36% 24.24% 34.20% 32.99% 22.11% 34.80%
Baseline (Curricularized) 21.93 58.27% 35.19% 25.43% 34.65% 33.09% 24.63% 35.25%
Layer Scaling Only 29.53 57.45% 32.58% 23.46% 30.00% 27.73% 22.62% 32.56%
MIDAS [39] 27.64 55.44% 32.58% 24.23% 29.62% 25.87% 23.81% 31.97%
CGLS ($N_{1}=6$) 26.58 61.70% 40.03% 23.55% 34.20% 28.92% 23.48% 36.28%
CGLS ($N_{1}=8$) 25.34 61.21% 41.37% 25.43% 34.65% 32.74% 26.55% 36.97%
Table 3: Evaluation results comparing CGLS and baseline approaches across average perplexity on The Pile dataset and core reasoning- and knowledge-intensive downstream tasks. CGLS, particularly the configuration starting with half of the final model depth, improves accuracy on downstream tasks like PIQA compared to the baselines. All models are compute-matched, but utilize much less compute than the pretrained variants of these models. Not compute-matched to the other methods, and trained on orders-of-magnitude more data. We provide pretrained results to illustrate a rough performance upper bound for the given architecture. The best model is bolded for each benchmark, with the second-best model underlined.
GPT-2-Small Scale.
The GPT-2 scale CGLS models achieve consistent improvements on downstream tasks (Table 3), and consistently lower perplexity across nearly all subsets of The Pile, as illustrated in Figure 2. For example, the $N_{1}=6$ model attains a 60.12% accuracy on PIQA, a 2.1-point improvement over the randomized baseline, along with gains on ARC-Challenge (22.53% vs. 21.08%).
LLaMA-3.2-1B Scale.
The larger models demonstrate clearer trends. While perplexity results diverged from the GPT-2-Small trend, with the randomized baseline achieving the lowest perplexity across all Pile domains (see A.1 for detailed perplexity results), CGLS nonetheless achieves the strongest downstream performance. CGLS with $N_{1}=8$ achieves the highest accuracy across nearly all downstream tasks, improving over the best baseline on PIQA (61.21% vs. 59.09%), ARC-Easy (41.37% vs. 36.36%), and MMLU-STEM (26.44% vs. 24.84%). The gains on ARC-Easy were especially striking, exceeding the best baseline by over 5 percent. Notably, the stacking baselines, without a synchronized data curricula, did not improve downstream performance. For MMLU, results remained near chance overall, reflecting the well-established difficulty of the benchmark for small models [32, 19]. However, token-normalized and PMI metrics revealed some structure. The STEM category, in particular, showed a meaningful increase: 26.44% for CGLS vs. 24.84% for the curricularized baseline and 21.92% for the randomized baseline. This suggests that CGLS may be particularly helpful in technical and reasoning-heavy domains. Further analysis is needed to understand rankings within MMLU subsets (see A.2 for further analysis), but the trend is promising.
@@ 4 Results
Comparing Progressive Scaling Strategies.
Among the GPT-2 CGLS variants, the best performance was consistently achieved with the $N_{1}=6$ model, starting from half the final model depth. This configuration outperformed both the ’Simple’ strategy and the $N_{1}=3$ model, suggesting that initializing from too small a model may hinder the transferability of learned representations. Similarly, for LLaMA-3.2-1B, CGLS with $N_{1}=8$ (half of the final 16-layer model) delivered the best downstream performance. Increasing the starting depth beyond half the final depth did not yield additional improvements (see A.1). This suggests that starting from half the model depth may strike a near-optimal balance between early-stage tractability and representational richness in progressive training frameworks.
5 Discussion
Inspired by how humans acquire cognitive capabilities in tandem with brain maturation and exposure to progressively more complex information, we align model capacity with data complexity during language model pretraining. We propose a new pretraining paradigm, Curriculum-Guided Layer Scaling (CGLS), that grows models in depth while adapting the distribution of training data to match their evolving capabilities. Applied to pretraining at GPT-2-Small and LLaMA-3.2-1B parameter counts, our approach yields consistent improvements in downstream task performance. In compute-matched comparisons, CGLS models outperform baselines on benchmarks requiring structured reasoning and domain-specific understanding, such as ARC, PIQA, and MMLU-STEM. Importantly, stacking-based baselines that expand model depth without synchronized data curricula do not improve downstream performance—CGLS significantly outperforms them across key benchmarks, demonstrating that aligning model growth with data complexity can unlock the benefits of progressive growing. Altogether, our results establish CGLS as a promising new paradigm for compute-efficient language model pretraining.
Limitations.
While CGLS yields improvements across several knowledge-intensive and reasoning tasks, in a few tasks such as LAMBADA and HellaSwag, CGLS models perform comparably to the baselines, suggesting that not all tasks benefit equally from the progressive curriculum. We also observe that CGLS does not always lower perplexity: at LLaMA-3.2-1B scale, the randomized baseline achieved slightly better perplexity despite underperforming on reasoning tasks. This disconnect between perplexity and functional generalization highlights the need for better evaluation frameworks.
Future Work.
We find in experiments that the initial model size is a crucial hyperparameter: starting with half the final number of layers (“$N_{1}=N/2$”) yielded the best performance at both pretraining scales. Starting too small (e.g., $N_{1}=3$ at GPT-2 scale or $N_{1}=6$ at Llama-3.2-1B scale) led to weaker final models, which we postulate may stem from insufficient capacity to master basic linguistic elements. Starting too large (e.g., $N_{1}=10$ for LLaMA) offered no additional benefit and, in some cases, underperformed standard training (results provided in A.1). Given our limited academic budget, we were not able to exhaustively tune these and other hyperparameters, suggesting that CGLS may be further improved. Future work could explore several directions: (1) more finely tuning the stage lengths and dataset proportions, (2) exploring other notions—e.g., syntactic diversity, lexical richness, or factuality—for defining the curriculum, (3) applying curriculum-aware optimization schedules (e.g., varying learning rates with data complexity), and (4) applying CGLS to non-text modalities (e.g., vision or multimodal data).
Broader Impacts.
CGLS is a general-purpose pretraining framework designed to improve the compute efficiency of language model pretraining. Potential positive impacts include democratizing access to high-performing language models, reducing pretraining compute for research groups with limited resources, and improving model interpretability through more structured growth. However, like other LLM training methods, CGLS could also be used to accelerate development of models deployed in harmful or unregulated ways. These include generating disinformation, bias amplification, or surveillance applications. While CGLS itself does not introduce new capabilities in these domains, it could reduce the cost and time needed to train capable models, slightly lowering the barrier for misuse. To mitigate risks, we recommend applying existing safety and alignment techniques during downstream fine-tuning stages, particularly for open-domain deployment. We also stress that our approach is most beneficial when paired with curated datasets and clear usage constraints.
Appendix A Technical Appendices and Supplementary Material
A.1 Additional Results
We present a more detailed comparison, including results for CGLS starting at 10 layers, in Table 4. CGLS with $N_{1}=8$ achieves the best overall downstream performance, including the highest accuracies on PIQA, ARC-E, ARC-Challenge, Hellaswag, and MMLU-STEM. CGLS with $N_{1}=6$ performs competitively, whilst starting with too large a model ($N_{1}=10$) underperforms on both perplexity and accuracy, supporting our hypothesis that synchronizing capacity with complexity is crucial.
Figure 3 further illustrates this trend, showing per-domain perplexity across The Pile. The randomized baseline achieves the lowest perplexity across all subsets, but CGLS outperforms other baselines such as MIDAS and layer-scaling without curricula, particularly on reasoning-heavy or technical subsets like Github and StackExchange.
Model
Pile
Perplexity
PIQA ARC-E ARC-C HellaSwag Lambada
MMLU
Llama-3.2-1B
Baseline (Randomized) 20.83 59.09% 36.36% 24.24% 34.20% 32.99% 22.11% 34.80%
CGLS ($N_{1}=6$) 26.58 61.70% 40.03% 23.55% 34.20% 28.92% 23.48% 36.28%
CGLS ($N_{1}=8$) 25.34 61.21% 41.37% 25.43% 34.65% 32.74% 26.55% 36.97%
CGLS ($N_{1}=10$) 31.69 57.34% 35.19% 23.38% 29.29% 24.96% 24.78% 32.49%
Table 4: Evaluation results comparing the compute-matched randomized baseline to CGLS across initial model sizes on The Pile perplexity and downstream tasks: PIQA, ARC-Easy, ARC-Challenge, Hellaswag, Lambada, and MMLU-STEM. The approach starting with half of the final model depth, achieves the highest accuracy among the CGLS experiments. The best model is bolded for each benchmark, with the second-best model underlined.
@@ A.1 Additional Results
Baseline
Randomized
Baseline
Curricularized
Layer
Scaling
MIDAS
CGLS
($N_{1}=6$)
CGLS
($N_{1}=8$)
Average 23.17% 24.93% 24.37% 23.50% 23.82% 25.10%
Other 24.35% 27.75% 27.18% 26.27% 27.63% 27.52%
Professional 24.20% 25.43% 25.66% 23.39% 25.38% 25.59%
Humanities 23.45% 24.72% 24.94% 24.56% 24.18% 23.99%
STEM 22.11% 24.63% 22.62% 23.81% 23.48% 26.55%
Social Sciences 21.98% 26.79% 23.90% 22.27% 23.07% 22.99%
biology 22.78% 22.78% 25.00% 22.92% 27.22% 27.78%
business 25.69% 25.21% 27.50% 22.92% 27.38% 26.30%
chemistry 18.81% 25.41% 22.44% 25.41% 20.46% 24.75%
computer science 26.21% 25.49% 23.54% 26.21% 23.06% 27.67%
economics 21.43% 27.09% 25.20% 20.89% 24.39% 25.07%
engineering 22.76% 22.76% 24.14% 25.52% 25.52% 26.21%
geography 18.69% 27.27% 18.69% 15.66% 18.69% 19.70%
health 23.05% 23.96% 26.89% 22.01% 26.22% 26.59%
history 23.44% 27.20% 24.73% 24.52% 23.87% 24.84%
law 24.56% 26.89% 23.65% 24.90% 23.65% 24.33%
math 20.49% 25.09% 20.11% 23.50% 21.43% 27.44%
other 24.35% 27.75% 27.18% 26.27% 27.63% 27.52%
philosophy 23.21% 23.71% 25.55% 24.65% 24.40% 23.56%
physics 22.81% 25.47% 23.28% 22.66% 23.91% 23.91%
politics 20.63% 25.46% 21.38% 21.75% 18.96% 19.52%
psychology 23.37% 27.10% 25.00% 24.30% 24.69% 23.76%
sociology 25.87% 23.38% 19.90% 23.88% 23.38% 24.38%
Table 5: MMLU accuracies across broader categories and finer subcategories for all Llama-3.2-1B scale models. We report zero-shot accuracy using PMI-based evaluation. The best-performing model for each category or subcategory is bolded, and the second-best is underlined. While the differences are often marginal, CGLS achieves the highest average accuracy and shows notable gains in STEM and reasoning-heavy subdomains.
A.2 MMLU Breakdown
Table 5 presents a finer-grained view of MMLU performance across individual categories and subcategories. While overall accuracy remains near chance, CGLS with $N_{1}=8$ achieves the highest average accuracy (25.10%), outperforming all baselines. The strongest gains appear in technical domains such as STEM (26.55%), math (27.44%), and computer science (27.67%), suggesting that progressive scaling may enhance performance in reasoning-heavy subfields. However, performance varies substantially across subcategories, with different models achieving the best result depending on the domain.
A.3 Experimental Setup
All training was performed using the Hugging Face Transformers library [46] on NVIDIA H100 80GB GPUs with NVLink. Each model was initialized with a learning rate of $2\times 10^{-4}$ and optimized using AdamW with a weight decay of 0.01. A warmup-stable-decay (WSD) learning rate schedule was applied, with 1,000 warmup steps during the initial stage (smallest model), followed by a stable learning rate across intermediate stages, and cosine decay to zero during the final stage (largest model). Layers added after model expansion were initially trained with a higher learning rate of $5\times 10^{-4}$, for 20% of the compute budget for the given stage. Optimizer states were reset between stages for all experiments involving layer expansion.
A.4 Dataset and Model Licenses
This work makes use of several publicly available datasets and pretrained models. The TinyStories dataset is released under the CDLA-Sharing-1.0 license and is composed of synthetic data generated for research purposes. The BookCorpus dataset, though widely used in academic NLP research, is no longer distributed and therefore does not have a formal license, though it may include copyrighted content. We include it here in line with standard usage in prior literature, but note its ambiguous legal status.
The DCLM (DataComp for Language Models) dataset is released under the Creative Commons Attribution 4.0 license as part of the DataComp benchmark suite, with a clear focus on research usability and content quality filtering.
For pretrained models, we use GPT-2-Small, which is released under the MIT model license. The LLaMA-3.2-1B model is made available by Meta under the LLaMA 3.2 Community License. While usable for research and certain commercial applications, the license prohibits redistribution of any model weights generated by Meta.

### v2

@@ Appendix A Appendix
A.1 Experimental Setup
All training was performed using the Hugging Face Transformers library (Wolf et al., 2019) on NVIDIA H100 80GB GPUs with NVLink. Each model was initialized with a learning rate of $2\times 10^{-4}$ and optimized using AdamW with a weight decay of 0.01. A warmup-stable-decay (WSD) learning rate schedule was applied, with 1,000 warmup steps during the initial stage (smallest model), followed by a stable learning rate across intermediate stages, and cosine decay to zero during the final stage (largest model). Layers added after model expansion were initially trained with a higher learning rate of $5\times 10^{-4}$, for 20% of the compute budget for the given stage. Optimizer states were reset between stages for all experiments involving layer expansion.
The full stage-wise splits of TinyStories, BookCorpus, and DCLM for the GPT-2-Small experiments are provided in Table 4. For Llama-3.2-1B, the stratified mixture weights over general, domain-specific, and specialized subsets of DCLM are in Table 5. For the domain shift experiment from DCLM (general web text) to StarCoder (Python code), we used a four-stage progression with mixture weights of 80/20, 70/30, 60/40, and 50/50 (DCLM/StarCoder), with all other hyperparameters identical to the other experiments.
Training a 1B parameter model with CGLS on 2.5B tokens took 5h53m on 4x H100s, while the compute-matched baseline took 5h39m on the same configuration. We note that our implementation of CGLS is not yet optimized for efficiency, e.g., with checkpoint loading from a fast disk. Moreover, no aspects of CGLS should meaningfully decrease model FLOPs utilization (MFUs), and CGLS and the baseline are compute-matched. Given this, we anticipate the wall-clock time of CGLS and the baseline should be essentially identical.
Model # Shots PIQA ARC-E ARC-C HellaSwag LAMBADA Average
@@ A.1 Experimental Setup
Baseline (Randomized) 0 54.41% 32.69% 23.98% 31.41% 29.89% 34.48%
5 56.31% 33.15% 26.19% 31.32% 31.32% 35.66%
Baseline (Curricularized) 0 53.86% 27.30% 30.19% 30.19% 34.02% 35.11%
5 54.73% 29.27% 29.73% 29.73% 33.57% 35.41%
Layer-Scaling Only 0 53.54% 26.71% 28.89% 28.46% 27.67% 33.05%
5 54.95% 26.71% 29.38% 29.11% 28.37% 33.70%
MIDAS 0 51.58% 26.54% 30.19% 28.89% 25.91% 32.62%
5 57.18% 28.33% 31.81% 29.38% 29.42% 35.22%
CGLS 0 56.69% 33.75% 29.18% 31.95% 31.04% 36.52%
5 58.65% 33.65% 29.10% 32.02% 31.95% 37.07%
Table 6: Zero- and 5-shot evaluation results across benchmarks for different training strategies. CGLS achieves the best average at both shot settings (bold).
Figure 2: Results for 1D hyperparameter sweeps over the starting model depth and stage-wise training budgets for the 1B-parameter model, trained on 2.5B tokens from DCLM. The average across the benchmarks is denoted with a dashed line.
A.2 Few Shot Results
In standard academic pretraining works, few-shot evaluations are uncommon because effective in-context learning emerges at larger scales, as shown in the GPT-3 paper Brown et al. (2020). For example, recent large-scale academic works such as H-Nets Hwang et al. (2025), which pretrains 1B parameter models for 100B tokens, do not include few-shot evaluations. Nevertheless, we conduct some few-shot evaluations, with PMI accuracy Gu et al. (2024) to reduce variance, and report the results in Table 6. For both zero- and 5-shot evaluation, CGLS achieves the best average across the benchmarks.
Model
PIQA
ARC-E ARC-C HellaSwag Lambada Average
@@ A.2 Few Shot Results
Baseline 59.18 $\pm$ 0.28% 36.46 $\pm$ 0.42% 22.38 $\pm$ 0.78% 27.81 $\pm$ 0.24% 20.49 $\pm$ 0.36% 33.26 $\pm$ 0.24%
CGLS 62.19 $\pm$ 1.42% 40.47 $\pm$ 2.43% 23.78 $\pm$ 1.76% 30.06 $\pm$ 1.24% 23.22 $\pm$ 1.19% 35.94 $\pm$ 1.60%
Table 7: Experimental replication results comparing CGLS and baseline approaches at the 2.5B token scale. Both models are compute-matched with training runs repeated three times in each case. We report zero-shot results with standard errors, with CGLS achieving a higher score across all benchmarks.
@@
Figure 5: Validation performance at each training stage for all methods. BL refers to the baseline, with R denoting the randomized variant, and C the curricularized variant.
A.3 Hyperparameter Ablations
We conduct ablations of core hyperparameters such as stage-wise training budgets and the starting depth. Though we are not able to exhaustively tune these hyperparameters on our academic compute budget, we report below some results conducting 1D sweeps on these key parameters. We find in figure 2 that starting at half the final depth ($N_{1}=8$) consistently yields the best results across downstream tasks, and that allocating 40% of the compute budget to the final stage performs best, outperforming front-heavy (25%) and back-heavy (50%, 75%) schedules. Increasing the starting depth beyond half the final depth did not yield additional improvements. These findings suggest that a balanced schedule that gradually expands capacity optimizes performance. Additionally, starting from half the model depth may strike a near-optimal balance between early-stage tractability and representational richness in progressive training frameworks.
A.4 Stability
Replication experiments at the 1B scale (Table 7) confirm that the improvements from CGLS are consistent across multiple seeds. While the variance across runs is slightly larger for CGLS than for the baseline, the mean accuracy is higher on every benchmark, indicating that the method’s benefits are robust. This suggests that progressive scaling may introduce some additional variability in optimization dynamics, but reliably yields stronger final models.
A.5 Perplexity Results
At the GPT-2-Small scale, CGLS achieves lower perplexity than the randomized and curricularized baselines on average, but the gains are most pronounced on challenging, knowledge-intensive subsets of The Pile such as PubMed Abstracts, PhilPapers, and Wikipedia (Figure 3).
For Llama-3.2-1B trained on 2.5B tokens, however, we observe the opposite trend: CGLS yields higher perplexity than the baselines across most subsets of The Pile (Figure 4). As discussed in the main paper, this divergence highlights that perplexity does not always align with downstream task improvements, particularly at smaller training budgets. Notably, scaling to a Chinchilla-optimal 20B tokens, this discrepancy disappears, and CGLS improves both perplexity and downstream accuracy.
A.6 Performance By Stage
Figure 5 reports the average accuracy across the benchmarks after each stage of training for CGLS and competing methods. The results highlight two main trends. First, CGLS improves steadily across stages, surpassing the randomized baseline by Stage 3 and continuing to gain performance in the final expansion. Second, while alternative progressive strategies such as MIDAS and LS also show early improvements, their performance plateaus sooner, converging below the baseline reference. In contrast, CGLS maintains upward momentum throughout the curriculum, suggesting that its synchronized data and model growth more effectively transfers knowledge from earlier stages into the final model.